Please answer 20 and 21, but you may need to use question 19. Thanks!

19. Colleen is a cash basis, calendar year individual taxpayer who has a filing status of single. Colleens taxable income20. (a) Given the same facts as in #19 above and assuming that Colleen wants to minimize her total tax liability, would she p
Answer #1

Answer to Question 20:

A. Coleen has a taxable income of $ 195000 in 2019 and the tax payable on her income works out to $ as worked out below

$ 157501 to $ 200000 : $ 32089.50

plus 32% of the amount over $157500 = 195000-157500= 37500 X 32% = $12000  

The total tax payable is therefore $ 32089.50 plus $12000 = $ 44089.50

Actual Tax Prepayment $ 52000   

If Coleen opts for 20% refundable tax credit she will get a refund of $ 52000 minus 44089.50 = $ 7910.50

But if she opts for adjusted gross income, she will get a refund of only the tax paid on the amount of deduction claimed i.e. $ 2000 which will be $ 640 (at her marginal rate of taxation of 32%).

Hence Coleen can minimize her tax liability by opting for 20% refundable tax credit.

20 (B)  Assuming the taxable income of Coleen at $24000, the net income of her in 2019 would be

Toatal Income: $24000

Less Standard deduction for single citizens: $12200

Net taxable income= $ 11800

The tax payable by her$2686 as below:

10% for income upto 9700 @10% = $970

9701 to 11800 @12% = 252 (11800-9700X0.12)

Total tax = 970+252 = 1222

In case Coleen opts for adjusted gross income, her adjusted total income would be $ 9800 as reduced by a further $ 2000 for childcare. In this case her total tax payable would be $970 + $20.4= 990.4. Hence, in this case her tax liability will be by reduce by opting for adjusted gross income which will reduce her tax liabitity by 1222-990.4 = $ 231.60.

Answer to Question 21:

21 A) The firm will report a taxable income of $ 3,00,000, as reduced by any tax rebates or concessions available to it as per the nature of its business (which is not clear from the question and hence, not possible to comment upon), . But as per taxation system for partnerships in USA, the tax will not be levied on the firm but its income will be added to the individual partners' income as per their respective shares in profits and taxed in their hands.

21 B) When a partner withdraws money from a partnership firm, he is presumed to wihdraw his share in profits first and thereafter any money he withdraws comes from his capital account. In the present case, Tom is entitled to 40% share in profits of the partnership firm. His share in the profits of the firm @40% is $ 120000. Hence, if he wthdraws only $ 90000, it will be only from profit and not capital. But, as per tax laws, Tom will still have to include his whole share in profits of the firm i.e. $ 120000 in his income for 2019 and pay tax on this amount.